# Szegő Theorem for Operator Orthogonal Polynomials

### 19 Mar 2026 v3

@@ 2 Preliminaries
Let $\mathcal{B}(\mathbb{T})$ denote the class of all Borel measurable subsets of the unit circle $\mathbb{T}$. We define $\mathcal{L}(\mathcal{H},\mathcal{K})$ as the algebra of all bounded linear operators from $\mathcal{H}$ to $\mathcal{K}$ where $\mathcal{H}$ and $\mathcal{K}$ are separable spaces. When $\mathcal{H}=\mathcal{K}$, we abbreviate this to $\mathcal{L}(\mathcal{H})$. Recall that $A\in\mathcal{L}(\mathcal{H})$ is positive definite $(A>0)$ if $\langle Ax,x\rangle>0$ for $0\neq x\in\mathcal{H}$, and $A$ has a bounded inverse.
@@ 2.3 Kolmogorov Isomorphism and Prediction Theory
where $I$ is the identity operator on $\mathcal{H}$, reformulates the Kolmogorov-Wiener prediction problem as an approximation problem in $L^{2}(\mu,\mathcal{H})$.
@@ 3 Operator orthogonal polynomials
here $C-B^{*}A^{-1}B>0$ is called the Schur complement of $A$.
@@ 6 Bernstein-Szegő Approximation
because the left-hand side is equal to the right-hand side for large enough $N$. Since Laurent polynomials are dense in $C(\mathbb{T})$, equation (27) holds for all $f$. In other words, we have weak convergence. ∎
@@ 7 The Szegő Limit Theorem
Before proving our results in the operator case, let us review some definitions related to infinite determinants [17]. For an operator $A$ that is trace class, the determinant $\mathrm{det}(I-A)$ is well-defined and can be written as
@@
where $(\lambda_{j})_{j=1}^{\infty}$ are the eigenvalues of $A$. On the other hand, if $A$ is a Hilbert-Schmidt operator, we define the second regularized determinant as
@@
This is based on the observation that $(I-A)e^{A}-I$ is trace class. It is worth noting that when $A$ is trace class, we have
@@
Moreover, for two Hilbert-Schmidt operators $A$ and $B$, we have

### 20 Mar 2026 v4

@@ 2 Preliminaries
Let $\mathcal{B}(\mathbb{T})$ denote the class of all Borel measurable subsets of the unit circle $\mathbb{T}$. We define $\mathcal{L}(\mathcal{H},\mathcal{K})$ as the algebra of all bounded linear operators from $\mathcal{H}$ to $\mathcal{K}$ where $\mathcal{H}$ and $\mathcal{K}$ are separable spaces. When $\mathcal{H}=\mathcal{K}$, we abbreviate this to $\mathcal{L}(\mathcal{H})$. Recall that $A\in\mathcal{L}(\mathcal{H})$ is positive definite $(A>0)$ if $\langle Ax,x\rangle>0$ for $0\neq x\in\mathcal{H}$, and AA has a bounded inverse.
@@ 2.3 Kolmogorov Isomorphism and Prediction Theory
where II is the identity operator on $\mathcal{H}$, reformulates the Kolmogorov-Wiener prediction problem as an approximation problem in $L^{2}(\mu,\mathcal{H})$.
@@ 3 Operator orthogonal polynomials
here $C-B^{*}A^{-1}B>0$ is called the Schur complement of AA.
@@ 6 Bernstein-Szegő Approximation
because the left-hand side is equal to the right-hand side for large enough NN. Since Laurent polynomials are dense in $C(\mathbb{T})$, equation (27) holds for all $f$. In other words, we have weak convergence. ∎
@@ 7 The Szegő Limit Theorem
Before proving our results in the operator case, let us review some definitions related to infinite determinants [17]. For an operator AA that is trace class, the determinant $\mathrm{det}(I-A)$ is well-defined and can be written as
@@
where $(\lambda_{j})_{j=1}^{\infty}$ are the eigenvalues of AA. On the other hand, if AA is a Hilbert-Schmidt operator, we define the second regularized determinant as
@@
This is based on the observation that $(I-A)e^{A}-I$ is trace class. It is worth noting that when AA is trace class, we have
@@
Moreover, for two Hilbert-Schmidt operators AA and $B$, we have